OpenStudy (anonymous):

Help! I give medals:) Pic below Solve for x in the right triangular prism.Show work

OpenStudy (anonymous):



OpenStudy (anonymous):

@nickersia do u think u could help me :)

OpenStudy (anonymous):

Surface area is (5*12)*2 + 5*x + 12*x + 13*x = 360 Can you work it out? :)

OpenStudy (anonymous):

Sorry (5*12)/2*2 So just 5*12 + 5*x + 12*x + 13*x = 360

OpenStudy (anonymous):

so solve the second one

OpenStudy (anonymous):

5*12 = 60 So, 5*x + 12*x + 13*x = 300 x(5+12+13)=300 x*30=300 x=10 :)

OpenStudy (anonymous):

Oh alright thanks for the help:)

OpenStudy (anonymous):

No problem :)